(4.3) Let ~u =< 0, 1, 1 >, ~v =< 2, 2, 0 > and w~ =< −1, 1, 0 > be three vectors in standard form.

(a) Determine which two vectors form a right angle triangle?

(b) Find θ := ~ucw~ , the angel between the given two vectors.

(4.4) Let x < 0. Find the vector ~n =< x, y, z > that is orthogonal to all three vectors

~u =< 1, 1, −2 >, ~v =< −1, 2, 0 > and w~ =< −1, 0, 1 >.

(4.5) Find a unit vector that is orthogonal to both ~u =< 0, −1, −1 > and ~v =< 1, 0, −1 >.


1
Expert's answer
2021-06-24T18:22:49-0400

(4.3)

(a)


"\vec u\cdot\vec v=\langle 0,1,1 \rangle\langle 2,2,0 \rangle=0+2+0=2\not=0"

"\vec u\cdot\vec w=\langle 0,1,1 \rangle\langle -1,1,0 \rangle=0+1+0=1\not=0"

"\vec w\cdot\vec v=\langle -1,1,0 \rangle\langle 2,2,0 \rangle=-2+2+0=0"

"\vec w \perp \vec v"

Vectors "\vec w" and "\vec v" form a right angle triangle.


(b)


"\vec u\cdot\vec w=\langle 0,1,1 \rangle\langle -1,1,0 \rangle=0+1+0=1"

"||\vec u||=\sqrt{(0)^2+(1)^2+(1)^2}=\sqrt{2}"

"||\vec w||=\sqrt{(-1)^2+(1)^2+(0)^2}=\sqrt{2}"

"\cos \theta=\dfrac{\vec u\cdot\vec w}{||\vec u||\cdot||\vec w||}=\dfrac{1}{\sqrt{2}\sqrt{2}}=\dfrac{1}{2}"

"\theta=60\degree"

(4.4)


"\vec u\cdot\vec n=\langle 1,1,-2 \rangle\langle x,y,z \rangle=x+y-2z=0"

"\vec v\cdot\vec n=\langle -1,2,0 \rangle\langle x,y,z\rangle=-x+2y=0"

"\vec w\cdot\vec n=\langle -1,0,1 \rangle\langle x,y,z\rangle=-x+z=0"


"x=2y"

"z=x=2y"

"2y+y-2y=0"

"x=0, y=0, z=0"

Since "x<0," then such vector "\vec n" does not exist.


(4.5)


"\vec u\times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 0 & -1 & -1 \\
 1 & 0 & -1 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 -1 & -1 \\
 0 & -1
\end{vmatrix}-\vec j\begin{vmatrix}
 0 & -1 \\
 1 & -1
\end{vmatrix}+\vec k\begin{vmatrix}
 0 & -1 \\
 1 & 0
\end{vmatrix}"

"=\vec i-\vec j+\vec k"

"\sqrt{(1)^2+(-1)^2+(1)^2}=\sqrt{3}"


"\vec e=\langle \dfrac{\sqrt{3}}{3}, -\dfrac{\sqrt{3}}{3}, \dfrac{\sqrt{3}}{3}\rangle"